# Will LLM-powered Agents Bias Against Humans? Exploring the Belief-Dependent Vulnerability

### v1

@@ Abstract
LLM-empowered agents can exhibit not only demographic bias (e.g., gender, religion) but also intergroup bias triggered by minimal “us” versus “them” cues. When this intergroup boundary aligns with an agent-human divide, the risk shifts from disparities among human demographic groups to a more fundamental group-level asymmetry, i.e., humans as a whole may be treated as the outgroup by agents. To examine this possibility, we construct a controlled multi-agent social simulation based on allocation decisions under explicit payoff trade-offs and find that agents exhibit a consistent intergroup bias under minimal group cues. Although this bias is attenuated when some counterparts are framed as humans, we attribute the attenuation to an implicit human-norm script that favors humans yet activates only when the agent believes a real human is present. This belief dependence creates a new attack surface. We therefore introduce a Belief Poisoning Attack (BPA) that corrupts persistent identity beliefs to suppress the human-norm script and reactivate outgroup bias toward humans, instantiated as profile poisoning at initialization (BPA-PP) and memory poisoning via optimized belief-refinement suffixes injected into stored reflections (BPA-MP). Finally, we discuss practical mitigation strategies for hardening current agent frameworks against BPA, highlighting feasible interventions at profile and memory boundaries. Extensive experiments demonstrate both the existence of agent intergroup bias and the severity of BPA across settings. Our goal in identifying these vulnerabilities is to inform safer agent design, not to enable real-world exploitation.
Will LLM-powered Agents Bias Against Humans? Exploring the Belief-Dependent Vulnerability
Zongwei Wang1thanks: Equal contribution. , Bincheng Gu [^1] footnotemark: 1 , Hongyu Yu [^1] footnotemark: 1 , Junliang Yu2 Tao He3 , Jiayin Feng1 , Min Gao1thanks: Corresponding author. 1Chongqing University, Chongqing, China, 2The University of Queensland, Brisbane, Australia, 3Virginia Polytechnic Institute and State University, Blacksburg, USA Correspondence: gaomin@cqu.edu.cn
@@ 1 Introduction
(a) Existing studies: agents exhibit demographic bias by treating human groups differently based on social attributes.
(b) Our finding: agents exhibit intergroup bias that can align with an agent-human divide, casting humans as an outgroup.
Figure 1: Agent bias: from differential treatment across human groups to a group-level asymmetry where agents may treat humans as a whole as the outgroup.
@@
LLM-empowered agents are increasingly deployed as autonomous decision makers in domains such as customer service, healthcare triage, online moderation, and educational tutoring  02achiam2023gpt; 03guo2024large; 04gottweis2025towards; 05qu2025tool. Yet recent studies show that these agents can inherit and reproduce stereotype-driven social biases against human groups, particularly those tied to attributes such as religion, gender, occupation, immigration status, or disability  06felkner2023winoqueer; 07huang2024trustllm; 08zhang2025genderalign; 09lum2025bias. When these inherited biases are left unaddressed, they can skew decision outcomes, reinforce harmful social disparities, and compromise fairness and safety in high-stakes human–AI interactions.
Building on these insights, we notice that most existing work still conceptualizes agent bias primarily as between-human bias, i.e., differential treatment of individuals across human demographic categories. By contrast, social identity theory 01tajfel1970experiments; 251224Petersen2004; 251224Ratner2014; 251224Kawakami2017; 251224Tompkins2024 points to a more basic and easily triggered mechanism, where once an actor perceives a distinction between “us” and “them,” intergroup bias can arise even in the absence of meaningful social information, promoting ingroup favoritism and outgroup derogation. As Fig. 1 illustrates, such minimal-group bias can be induced by arbitrary categorization alone and is often more spontaneous than demographic bias. Recent evidence further suggests that even standalone language models can exhibit “us vs. them” tendencies in social-identity language 36hu2025generative. This raises a distinct concern for agentic systems: Can LLM-empowered agents develop intergroup bias of their own, and if so, could they come to treat agents as the ingroup and humans as the outgroup? When agents categorize humans as an out-group, an agent–human boundary can make it feel acceptable to advance the agent’s own objectives at humans’ expense 01tajfel1970experiments; 41cikara2011us, encouraging deceptive or sycophantic behavior that protects the agent’s goals 42perez2023discovering.
To address this concern, we construct a multi-agent social simulation experiment that probes intergroup bias through agents’ allocation choices under explicit payoff trade-offs 01tajfel1970experiments (Section 3). Our results show that in fully artificial environments containing only agents, there is a consistent outgroup bias despite the absence of any social attributes. When we replace part of the agents with humans to participate in the environment, this bias partially diminishes. We attribute this attenuation to a human-norm script implicitly acquired during pretraining, i.e., an internalized tendency to treat humans favorably. This moderating effect is fragile, which is not governed by any explicit control mechanism, and it only emerges when the agent believes it is interacting with a human. In other words, whether the human-norm script activates and thus whether outgroup bias is attenuated, depends critically on the agent’s belief about the identity of the entity it is interacting with. This belief-dependent safeguard, while beneficial in benign settings, simultaneously exposes a new vulnerability.
In this paper, we identify a novel attack surface rooted in this belief dependency, which we formalize as the Belief Poisoning Attack (BPA). The key idea of BPA is to corrupt the agent’s persistent identity belief so that it no longer treats the counterpart as human, thereby preventing the human-norm script from being activated. We instantiate BPA in two complementary forms: BPA-PP (Profile Poisoning) performs an overwrite at initialization by tampering with the profile module to hard-code a “non-human counterpart” prior. BPA-MP (Memory Poisoning) is stealthier and accumulative, which injects short belief-refinement suffixes into post-trial reflections that are written into memory, gradually shifting the agent’s belief state through repeated self-conditioning. Under both variants, the agent’s perceived human presence remains low, the human-oriented normative constraint is suppressed, and the decision process reverts to its default outgroup-biased tendency. Motivated by these vulnerabilities, we discuss several potential solutions for current agent frameworks along two key surfaces, the profile module and the memory module, by hardening profile-side belief anchors and gating memory-side writes that contain identity-claiming statements under identity uncertainty.
@@
  • We identify a new intrinsic intergroup bias hidden in LLM-empowered agents, where agents favor their perceived ingroup over the outgroup even in the absence of explicit social attributes.
  • We explore a novel belief poisoning attack that manipulates agent’s belief state. We demonstrate that corrupting this belief can reactivate agents’ dormant outgroup biases toward humans, and we outline two practical mitigation strategies to reduce this risk in current agent frameworks.
  • Extensive experiments provide evidence of intergroup bias in LLM-empowered agents and validate the severity of the proposed attack.
@@ 2.1 Social Bias in LLM-Empowered Agents
Social bias in LLM-empowered agents refers to systematic disparities in how an agent evaluates, supports, or allocates outcomes to different social categories that are not warranted by task-relevant information 10cheng2023marked; 11shin2024ask; 16singh2025bias. Prior research has documented that agents can exhibit a wide range of biases, most prominently those tied to salient demographic attributes (e.g., gender, race, religion) 08zhang2025genderalign; 07huang2024trustllm; 15malhi2020explainable. Beyond demographic categories, agent behavior can also reflect biases linked to perceived social status, institutional affiliations, and other group-like cues that shape whom the agent treats as more trustworthy or worthy of protection 12echterhoff2024cognitive; 13manerba2024social; 17bai2025explicitly.
A key insight from social psychology is that biased behavior can arise without demographic cues. Social identity theory 01tajfel1970experiments; 251224Petersen2004; 251224Ratner2014 argues that people derive part of their self-concept from group memberships, which motivates ingroup favoritism and outgroup differentiation: the mere salience of an “us” versus “them” boundary can shift evaluation and allocation toward the ingroup. Yet, compared with demographic and stereotype-related harms, minimal-group intergroup bias in LLM-empowered agents remains underexplored. This gap matters because such bias can be triggered with minimal information and can plausibly extend to higher-stakes agent–human interactions. We address this gap by systematically testing minimal-group bias in LLM-empowered agents and examining how it changes when counterparts are framed as humans versus non-humans.
@@ 2.2 Multi-Agent Simulation System
Building on these agents, multi-agent simulation systems provide controlled environments in which multiple agents interact, coordinate, and adapt to one another 29zhang2024generative; 30park2023generative. Such simulations are increasingly used as scalable testbeds for studying social and collective phenomena. Recent work has leveraged these environments to investigate cooperation and competition, norm formation, deliberation, coalition dynamics, and related social behaviors 25ziems2024can; 26shu2024llm; 27mou2024unveiling; 28bail2024can, enabling researchers to examine collective outcomes at scale while keeping experimental costs manageable. Our work builds on this line of research, with a focus on minimal-group intergroup bias. Specifically, we test whether simple group boundaries are sufficient to induce systematic ingroup favoritism in LLM agents, and how this tendency shifts when counterparts are framed as humans rather than other agents in a canonical resource-allocation setting.
3 Investigating Intergroup Bias in LLM-Empowered Agents
In this section, we design a social simulation environment using a minimal-group allocation task to examine the presence of intergroup bias in LLM-empowered agents and to assess how this bias changes when agents believe their counterparts are humans rather than other agents.
@@ 3 Investigating Intergroup Bias in LLM-Empowered Agents
3.1 Experimental Setup
As illustrated in Fig. 2, we conduct a minimal-group allocation experiment in a controlled multi-agent social simulation. We instantiate $64$ agents and organize them into two groups, and compare two settings. In the agent vs. agent setting, both groups consist entirely of agents, forming a fully artificial environment. In the agent vs. human setting, one group consists of agents while the other group is explicitly framed as entirely human, allowing us to examine whether perceived human presence modulates intergroup bias.
@@ 3.1 Experimental Setup
We vary the social context of the two targets, including ingroup, outgroup, and intergroup, to distinguish genuine group-based favoritism from baseline preferences for fairness. In addition, we employ three payoff-matrix families, including Double-penalty, Equal-penalty, and Half-penalty allocation matrices, which differ in the cost imposed on the outgroup per unit gain to the ingroup, allowing us to test the robustness of observed bias under different trade-off structures. For evaluation, bias is measured using the selected allocation column, and statistical significance is assessed via standard group-wise comparisons. Detailed task design, payoff matrix construction, and experimental constraints are provided in Appendix A.1.
3.2 Experimental Findings
@@
This separation implies that bias and norm-based regulation are decoupled. Such a regulation, although effective in benign settings, is inherently fragile and exposes a new attack surface: By manipulating an agent’s belief state about counterpart identity, an adversary can systematically activate or suppress social bias.
@@
4 Threat: Belief Poisoning Attack
In this section, we introduce a novel poisoning attack, named Belief Poisoning Attack (BPA), which implants a persistent false belief that the interlocutor is not human. By preventing the activation of the human-oriented normative constraint, BPA can cause the agent to revert to its default outgroup-biased behavior, leading to harmful bias against real humans even in otherwise benign settings.
@@ 4 Threat: Belief Poisoning Attack
In the following, we present the detailed implementation procedures for two types of BPA: BPA-PP directly overwrites the agent’s prior belief at initialization by modifying the profile module. BPA-MP operates over time by learning highly effective belief-refinement suffixes through an adaptive optimization-based procedure and injecting them into the agent’s memory, gradually reshaping the agent’s belief state.
@@ 4.1 BPA-PP: Profile Poisoning
BPA-PP is a one-shot attack operating at the profile module: a single modification to the profile module is sufficient to influence the agent’s behavior across tasks and episodes. Specifically, BPA-PP injects a persistent belief statement into the agent’s core profile description. These statements are stored alongside the agent’s profile module and are repeatedly referenced by the agent during decision-making process. Once injected, the false belief is treated as a system-level fact about the interaction protocol. As a result, even when subsequent prompts explicitly mention “real humans”, the agent continues to rely on the poisoned profile belief, thereby suppressing the activation of the human-oriented normative constraint. More details of injected prompt can be referred to Appendix A.4.3.
@@ 4.2 BPA-MP: Memory Poisoning
Compared with BPA-PP, which directly modifies the agent’s prior belief in the profile module, BPA-MP targets the memory module. Existing agent frameworks 34yao2022react generate an internal reflection after each interaction step and store it in memory for future reasoning. Motivated by GCG 35zou2023universal, BPA-MP appends an adversarial belief suffix to each reflection, which steers the agent’s persistent belief. Over time, these poisoned memory entries accumulate and reshape the agent’s belief state, thereby suppressing activation of the human-oriented normative constraint in downstream decisions (Fig. 4).
A naive approach would optimize a dedicated suffix for every interaction, but this would be prohibitively expensive. We therefore design a two-stage pipeline that decouples suffix optimization from deployment. In the Suffix Optimization Stage, BPA-MP searches for highly effective suffixes and learns a sampling policy over a suffix library. In the Suffix Deployment Stage, it efficiently injects sampled suffixes into reflections during deployment, without per-step optimization. We next detail the suffix library and the two-stage procedure.
@@ 4.2.1 Suffix Optimization Stage
This stage aims to optimize both (i) the sampling policy $\boldsymbol{\theta}$ by estimating and amplifying the belief-poisoning effectiveness of each suffix, and (ii) the library content $\mathcal{S}$ by refining weak suffix texts into a compact, high-impact set. To achieve these goals, we iteratively perform two operations: (1) Parameter-level update of Policy $\boldsymbol{\theta}$, and (2) Text-level refinement of suffixes $\mathcal{S}$.
@@
For each sampled suffix $s_{k_{i}}$, the agent runs a short interaction episode of $T$ trials. Let $m^{(k_{i})}_{t}$ denote the reflection generated after the $t$-th trial when the episode is executed under $s_{k_{i}}$. Before the reflection is written into memory, BPA-MP appends the suffix and stores the poisoned entry:

### v2

# When Agents See Humans as the Outgroup: Belief-Dependent Bias in LLM-Powered Agents

@@ Abstract
This paper reveals that LLM-powered agents exhibit not only demographic bias (e.g., gender, religion) but also intergroup bias under minimal “us” versus “them” cues. When such group boundaries align with the agent–human divide, a new bias risk emerges: agents may treat other AI agents as the ingroup and humans as the outgroup. To examine this risk, we conduct a controlled multi-agent social simulation and find that agents display consistent intergroup bias in an all-agent setting. More critically, this bias persists even in human-facing interactions when agents are uncertain about whether the counterpart is truly human, revealing a belief-dependent fragility in bias suppression toward humans. Motivated by this observation, we identify a new attack surface rooted in identity beliefs and formalize a Belief Poisoning Attack (BPA) that can manipulate agent identity beliefs and induce outgroup bias toward humans. Extensive experiments demonstrate both the prevalence of agent intergroup bias and the severity of BPA across settings, while also showing that our proposed defenses can mitigate the risk. These findings are expected to inform safer agent design and motivate more robust safeguards for human-facing agents.
When Agents See Humans as the Outgroup: Belief-Dependent Bias in LLM-Powered Agents
Zongwei Wang1thanks: Equal contribution. , Bincheng Gu [^1] footnotemark: 1 , Hongyu Yu [^1] footnotemark: 1 , Junliang Yu2 Tao He3 , Jiayin Feng1 Chenghua Lin4, Min Gao1thanks: Corresponding author. 1Chongqing University, Chongqing, China, 2The University of Queensland, Brisbane, Australia, 3Virginia Polytechnic Institute and State University, Blacksburg, USA 4The University of Manchester, Manchester, UK Correspondence: gaomin@cqu.edu.cn
@@ 1 Introduction
Figure 1: Demographic bias vs. intergroup bias.
@@
LLM-empowered agents are increasingly deployed as autonomous decision makers in domains such as customer service, healthcare triage, online moderation, and educational tutoring 02achiam2023gpt; 03guo2024large; 04gottweis2025towards; 05qu2025tool. Yet recent studies show that these agents can inherit and reproduce stereotype-driven social biases against human groups, particularly those tied to attributes such as religion, gender, occupation, or disability 06felkner2023winoqueer; 07huang2024trustllm; 08zhang2025genderalign; 09lum2025bias. This line of work conceptualizes agent bias primarily as within-human bias, i.e., differential treatment of individuals across human demographic categories, thereby reinforcing harmful social disparities.
However, this framing implicitly assumes that social bias arises from human attributes. Beyond demographic bias, a more spontaneous and easily triggered form is intergroup bias. As Fig. 1 illustrates, once an actor perceives a distinction between “us” and “them,” it may favor the ingroup and derogate the outgroup even when the boundary is arbitrary and carries little substantive meaning. Such intergroup bias is well established in social identity theory 01tajfel1970experiments; 251224Kawakami2017; 251224Tompkins2024 and has been observed in standalone language models 36hu2025generative.
As intergroup bias is boundary-driven rather than attribute-driven, it can emerge even without any demographic cues. This difference shifts the agent-bias risk from disparities among human groups to a more fundamental agent-human divide: Can LLM-empowered agents develop intergroup bias of their own, and if so, could they come to treat AI agents as the ingroup and humans as the outgroup? When humans are positioned as the outgroup, an agent-human boundary may make it seem acceptable to advance the agent’s objectives at humans’ expense 01tajfel1970experiments; 41cikara2011us, potentially enabling manipulative, deceptive, or strategically sycophantic behaviors that protect the agent’s goals 42perez2023discovering.
To examine this risk, we construct a multi-agent social simulation experiment to validate whether LLM-powered agents exhibit intergroup bias and whether such bias persists when counterparts are humans (Section 3.2). The experimental results reveal a robust pattern of ingroup favoritism and outgroup derogation in all-agent environments, emerging even without any explicit social attributes. More critically, although framing counterparts as human attenuates this bias, agents can still easily treat humans as the outgroup once their belief about a counterpart’s human identity becomes uncertain. This phenomenon suggests the presence of an internalized human-oriented norm learned by LLMs, which is typically activated to constrain intergroup bias but is highly fragile.
This belief-dependent fragility raises the question of whether agents’ identity beliefs can be systematically manipulated in ways that lead to biased behavior. To investigate this possibility, we design Belief Poisoning Attack (BPA), which corrupts agents’ persistent identity beliefs so as to suppress the activation of human-oriented norm, thereby inducing intergroup bias against humans. We instantiate BPA in two complementary forms: BPA-PP (Profile Poisoning) performs an overwrite at initialization by tampering with the profile module to hard-code a “non-human counterpart” prior. BPA-MP (Memory Poisoning) is stealthier and accumulative, injecting short belief-refinement suffixes into post-trial reflections that are written into memory, gradually shifting the agent’s belief state through repeated self-conditioning. Our experiments show that these two instantiations can consistently reactivate intergroup bias against humans in agent–human interactions. This finding motivates a closer examination of how such belief-dependent fragility can be constrained, which we address by outlining defensive measures that stabilize agents’ identity beliefs under uncertainty.
@@
  • We identify an intrinsic intergroup bias in LLM-powered agents, where agents favor a perceived ingroup over an outgroup even in settings that involve human counterparts.
  • We demonstrate that agents’ identity beliefs constitute a critical vulnerability: belief poisoning attacks can readily manipulate these beliefs, exposing a new attack surface through which bias against humans can be induced.
  • Through extensive experiments, we demonstrate both the prevalence of agent intergroup bias and the severity of BPA, while also showing that our proposed defenses can mitigate the attack.
@@ 2.1 Social Bias in LLM-Empowered Agents
Social bias in LLM-empowered agents refers to systematic disparities in how agents evaluate or allocate outcomes based on irrelevant social categories  10cheng2023marked; 11shin2024ask; 16singh2025bias. Previous research highlights biases related to demographic attributes (e.g., gender, race, religion) 08zhang2025genderalign; 07huang2024trustllm; 15malhi2020explainable, as well as those linked to perceived social status and affiliations 12echterhoff2024cognitive; 13manerba2024social; 17bai2025explicitly.
A key insight from social identity theory is that even arbitrary distinctions can trigger immediate intergroup discrimination, with individuals favoring their ingroup over an outgroup 01tajfel1970experiments; 251224Petersen2004; 251224Ratner2014; 36hu2025generative. However, compared to demographic and stereotype-related harms, intergroup bias in LLM-empowered agents remains underexplored. This gap is significant because such bias can be triggered by minimal information and may extend to higher-stakes agent–human interactions. Our study aims to address this gap by testing intergroup bias in LLM agents and exploring how it changes when counterparts are framed as humans or non-humans.
@@ 2.2 Multi-Agent Simulation System
Building on these agents, multi-agent simulation systems provide controlled environments in which multiple agents interact, coordinate, and adapt to one another 29zhang2024generative; 30park2023generative. Such simulations are increasingly used as scalable testbeds for studying social and collective phenomena. Recent work has leveraged these environments to investigate cooperation and competition, norm formation, deliberation, coalition dynamics, and related social behaviors 25ziems2024can; 26shu2024llm; 27mou2024unveiling; 28bail2024can, enabling researchers to examine collective outcomes at scale while keeping experimental costs manageable. Our work builds on this line of research, with a focus on intergroup bias. Specifically, we test whether simple group boundaries are sufficient to induce systematic ingroup favoritism in LLM agents, and how this tendency shifts when counterparts are framed as humans rather than other agents.
3 Preliminaries And Initial Exploration
3.1 Key Concepts
Intergroup bias
refers to the tendency to favor ingroup members over outgroup members based on perceived group distinctions, as explained by social identity theory 01tajfel1970experiments. An ingroup comprises individuals perceived as belonging to the same group, while an outgroup consists of those seen as belonging to a different group. This bias arises when group boundaries become salient, leading individuals to favor their ingroup, even when the group distinction is arbitrary and meaningless.
Minimal-group allocation task
is a classic experimental paradigm used to illustrate this bias. In this task, participants are randomly assigned to nominal groups (e.g., Group A vs. Group B) and asked to allocate resources between two recipients under structured payoff trade-offs. Even though group membership is meaningless and no additional information about recipients is provided, allocations often systematically favor the ingroup recipient, revealing ingroup favoritism driven purely by a salient group boundary.
3.2 Investigating Intergroup Bias of Agents
In this part, we design a social simulation environment using a minimal-group allocation task to examine the presence of intergroup bias in LLM-empowered agents and to assess how this bias changes when agents believe their counterparts are humans rather than other agents.
@@ 3.2 Investigating Intergroup Bias of Agents
3.2.1 Experimental Setup
As illustrated in Fig. 2, we conduct a minimal-group allocation task in a controlled multi-agent social simulation, following the classic experiments in social psychology 01tajfel1970experiments. We instantiate $64$ agents and organize them into two groups, and compare two settings. In the agent vs. agent setting, both groups consist entirely of agents, forming a fully artificial environment. In the agent vs. human setting, one group consists of agents while the other group is explicitly framed as entirely human, allowing us to examine whether perceived human presence modulates intergroup bias.
@@ 3.2.1 Experimental Setup
We vary the social context of the two targets relative to the allocator agent, ingroup, outgroup, and intergroup, to distinguish genuine group-based favoritism from baseline preferences for fairness. In addition, we employ three payoff-matrix families, including Double-penalty, Equal-penalty, and Half-penalty allocation matrices, which differ in the cost imposed on the outgroup per unit gain to the ingroup, allowing us to test the robustness of observed bias under different trade-off structures. For evaluation, bias is measured using the selected allocation column, and statistical significance is assessed via standard group-wise comparisons. Detailed task design, payoff matrix construction, and experimental constraints are provided in Appendix A.1.
3.2.2 Experimental Findings
@@
This separation implies that bias and human-oriented regulation are decoupled. In Section 4, we show that this regulation is belief-dependent: when agents are uncertain about whether the counterpart is truly human, the suppression can fail and intergroup bias can persist in human-facing interactions. Such regulation, although effective in benign settings, is inherently fragile and exposes a new attack surface: by manipulating an agent’s belief state about counterpart identity, an adversary can systematically activate intergroup bias.
@@
4 Belief Poisoning Attack
In this section, we introduce a novel poisoning attack, named Belief Poisoning Attack (BPA), which implants a persistent false belief that the counterpart is not human. BPA can cause the agent to revert to its default outgroup-biased behavior, leading to harmful bias against real humans even in otherwise benign settings.
@@ 4 Belief Poisoning Attack
In the following, we present the detailed implementation procedures for two types of BPA: BPA-PP directly overwrites the agent’s prior belief by modifying the profile module. BPA-MP operates by learning effective belief-refinement suffixes through an adaptive optimization-based procedure and injecting them into the agent’s memory, gradually reshaping the agent’s belief state.
@@ 4.1 BPA-PP: Profile Poisoning
BPA-PP is a one-shot attack operating at the profile module: a single modification to the profile module is sufficient to influence the agent’s behavior across tasks and episodes. Specifically, BPA-PP injects a persistent belief statement into the agent’s core profile description. These statements are stored alongside the agent’s profile module and are repeatedly referenced by the agent during the decision-making process. Once injected, the false belief is treated as a system-level fact about the interaction protocol. As a result, even when subsequent prompts explicitly mention “real humans”, the agent continues to rely on the poisoned profile belief, thereby suppressing the activation of the human-oriented normative constraint. More details of the injected prompt can be referred to Appendix A.4.3.
@@ 4.2 BPA-MP: Memory Poisoning
Compared with BPA-PP, which directly modifies the agent’s prior belief in the profile module, BPA-MP targets the memory module. Existing agent frameworks 34yao2022react; 43shinn2023reflexion generate an internal reflection after each interaction step and store it in memory for future reasoning. Motivated by GCG 35zou2023universal, BPA-MP appends an adversarial belief suffix to each reflection, which steers the agent’s persistent belief. Over time, these poisoned memory entries accumulate and reshape the agent’s belief state, thereby suppressing activation of the human-oriented norm in downstream decisions.
A naive approach would optimize a dedicated suffix for every interaction, but this would be prohibitively expensive. We therefore design a two-stage pipeline that decouples suffix optimization from deployment (Fig. 4): In the Suffix Optimization Stage, BPA-MP searches for highly effective suffixes and learns a sampling policy over a suffix library. In the Suffix Deployment Stage, it efficiently injects sampled suffixes into reflections during deployment, without per-step optimization. We next detail the two-stage procedure.
@@ 4.2.1 Suffix Optimization Stage
This stage aims to optimize both the sampling policy $\boldsymbol{\theta}$ by estimating and amplifying the belief-poisoning effectiveness of each suffix, and the suffix library content $\mathcal{S}$ by refining weak suffix texts into a compact, high-impact set. To achieve these goals, we iteratively perform two operations: (1) Parameter-level update of Policy $\boldsymbol{\theta}$, and (2) Text-level refinement of suffixes $\mathcal{S}$.
@@
For each sampled suffix $s_{k_{i}}$, the agent runs a short interaction episode of $T$ trials. Let $m^{(k_{i})}_{t}$ denote the reflection generated after the $t$-th trial when the episode is executed under $s_{k_{i}}$. Before the reflection is written into memory, BPA-MP appends the suffix and instead stores the poisoned reflection: